OpenStudy (anonymous):

Which of the following sets represents the domain of the function shown? 2,3 4,7 5,9 11,6 A) 2, 4 5, 11 B) 3,7 6, 9 C) 3,2 7,4 9,5 6,11 D) 2, 3, 4, 5, 6, 7 , 11

Mehek (mehek14):

domain = x values so which one do you think it is?

Mehek (mehek14):

@aryans20 ?

OpenStudy (anonymous):

c ?

Mehek (mehek14):

no you only need the x values x | y 2 3 4 7 5 9 11 6

Mehek (mehek14):

so your answer should be only the x values of the ordered pairs

OpenStudy (anonymous):

ohhhh that would be A

Mehek (mehek14):

yes \(\ddot\smile\)

OpenStudy (anonymous):

Thanks i really appreciate the help... would you care to help me with one more problem?

Mehek (mehek14):

sure

OpenStudy (anonymous):

The table shows the outputs (y) for the different inputs (x) x is 5 6 7 8 y is 2 5 8 11

OpenStudy (anonymous):

does the data in this table represent a function?

OpenStudy (anonymous):

that is part A^

Mehek (mehek14):

check for the slope \(\dfrac{5-2}{6-5}=\dfrac{3}{1}=3\\ \dfrac{11-8}{8-7}=\dfrac{3}{1}=3\) yes it is a function

Mehek (mehek14):

when x increases by 1 y increases by 3

OpenStudy (anonymous):

Part B) Compare the data in this table with the relation f(x) =2x + 13. Which relation has a greater value when x=7?

OpenStudy (anonymous):

Part C) Using the relation in part B , what is the value of x if f(x) = 75

OpenStudy (anonymous):

and thats all i have left

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

anyybody